(3.1) Find the point of intersection between the lines: < 3, −1, 2 > + t < 1, 1, −1 > and (3) < −8, 2, 0 > + t < −3, 2, −7 >.

(3.2) Show that the lines x + 1 = 3t, y = 1, z + 5 = 2t for t ∈ R and x + 2 = s, y − 3 = −5s, (5) z + 4 = −2s for t ∈ R intersect, and find the point of intersection.

(3.3) Find the point of intersection between the planes: −5x + y −2z = 3 and 2x −3y + 5z = −7


1
Expert's answer
2021-07-05T16:52:13-0400

(3.1)


"3+s=-8-3t""-1+s=2+2t""2-s=0-7t"


"s=-3t-11""4=-5t-10""1=-5t+2"



"s=-3t-11""3=-12""1=-5t+2"

No solution.

There is no point of intersection.


(3.2)


"3t-1=s-2""1=-5s+3""2t-5=-2s-4"


"s=3t+1""5s=2""2t=-2s+1"



"t=-\dfrac{1}{5}""s=\dfrac{2}{5}""t=\dfrac{1}{10}"


No solution.

There is no point of intersection.


(3.3)


"-5x+y-2z=3""2x-3y+5z=-7"



"-13x-z=2""2x-3y+5z=-7"

"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{2}{3}x+\dfrac{5}{3}z+\dfrac{7}{3}"



"x=-\dfrac{2}{13}-\dfrac{1}{13}z"

"y=\dfrac{29}{13}+\dfrac{21}{13}z"

"z\in\R"

The intersection between the planes is the line :<3, - 1,2> +<1, 1, - 1>


"\langle-\dfrac{2}{13}, \dfrac{29}{13}, 0\rangle+t\langle-\dfrac{1}{13}, \dfrac{1}{13},1\rangle, t\in \R"